OpenStudy (anonymous):

Which percent is equivalent to 2 and 5/6

OpenStudy (anonymous):

@amistre64 easy one for you

OpenStudy (anonymous):

@Firejay5

OpenStudy (anonymous):

np

OpenStudy (amistre64):

how do we find a percent?

OpenStudy (amistre64):

take the decimal version and multiply by 100 or the mixed numeral version if it simplifies good enough

OpenStudy (firejay5):

@jobindog 2 * 100 = 200% 5/6 = 0.8333333333333333 * 100 = 83.33%